OpenStudy (anonymous):

Simplify: 4√2 + √32 - √16

OpenStudy (luigi0210):

factor time!

OpenStudy (luigi0210):

\[4\sqrt{2}+\sqrt{4*4*2}-\sqrt{4*4}\]

OpenStudy (anonymous):

huh?